# Increasing Trust in Language Models through the Reuse of Verified Circuits

### v8

@@ Abstract
Language Models (LMs) are increasingly used for a wide range of prediction tasks, but their training can often neglect rare edge cases, reducing their reliability. Here, we define a stringent standard of trustworthiness whereby the task algorithm and circuit implementation must be verified, accounting for edge cases, with no known failure modes. We show that a model can be trained to meet this standard if built using mathematically and logically specified frameworks. In this paper, we fully verify an auto-regressive transformer model that performs n-digit integer addition. To exhibit the reusability of verified modules, we insert the trained integer addition model into a larger untrained model and train the combined model to perform both addition and subtraction. We find extensive reuse of the addition circuits for both tasks, easing verification of the more complex subtractor model. We discuss how inserting verified task modules into LMs can leverage model reuse to improve verifiability and trustworthiness of LMs built using them. The reuse of verified circuits reduces the effort to verify more complex composite models which we believe to be a significant step towards safety and interpretability of LMs.
Increasing Trust in Language Models through the Reuse of Verified Circuits
Philip Quirkethanks:    Core research contributor., Clement Neo†‡, Fazl Barezfootnotemark: †‡ Apart Research, University of Oxford
@@ 1 Introduction
Figure 1: An overview of our methodology: (1) We trained an accurate 6-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
@@
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need for understanding their behavior. Hendrycks and Mazeika (2022) argue that making these models interpretable is key to their safe use. Mechanistic interpretability focuses on demystifying and validating the algorithms behind model weights, translating complex computations into more human-understandable components (Raukur et al., 2022). This understanding aids in predicting model behavior in new situations and fixing model errors.
In creating and training a model, we aim for high accuracy and trustworthiness. We achieve this by holding the model to a standard we term known-good. We define a model performing a task to be known-good if:
  1. 1.
    The model’s algorithm for the task and the mechanisms it implements (the “circuits") are understood.
  2. 2.
    All possible task edge cases have been identified and tested.
  3. 3.
    Empirically, the model prediction accuracy is 99.9999% (a standard reliability measure used in industry and abbreviated as “six nines"). That is, it can perform the task one million times with at most 1 wrong prediction.
Exhaustive testing of a model task may be infeasible. For instance, when adding two 5-digit integers (e.g. 12345+67890) there are ten billion variations. Some tasks can be conceptualized within an existing formal framework that allows identification of all edge cases the model must handle. For example, in 5-digit addition, the most uncommon edge case is 55555+44445=100000, which requires a carry bit to cascade through all digits, occurring in only 0.002% of cases. A known-good model must incorporate algorithms to manage all known edge cases. A known-good model must have verified [^1] In this paper, ’verified’ has the mechanistic interpretability meaning that a specific group of interconnected neurons within a neural network reliably and causally contributes to the model’s output in a meaningful, understandable way, with supporting empirical evidence. circuits that perform the task accurately.
In this paper, we detail the development and interpretation of a known-good model for addition. Our findings indicate that the model constructs a specific circuit for each edge case, with these circuits sharing intermediate results. We confirm the validity of the entire set of circuits, ensuring they cover all identified edge cases. The model achieves a very low training loss and has six nines (99.9999%) accuracy. The model hence achieves our known-good standard. Additionally, we develop a “mixed" model capable of both addition and subtraction, incorporating the known-good addition model. This mixed model has six nines accuracy, and extensively reuses the addition circuits for both operations, facilitating the interpretation of the model’s algorithm. We make progress toward a known-good model for both addition and subtraction.
Hence, our main contributions are three-fold:
  • Defining several known-good n-digit addition models with six nines accuracy which all use the same algorithm.
  • Demonstrating a proof of concept for re-using a known good model in the training of another larger, more-capable model, simplifying the interpretation of the new model’s algorithm.
  • Defining several n-digit addition and subtraction models with six nines accuracy, that reuse established addition model circuits for both operations, and detailing progress towards these models being known-good.
@@ 2 Related Work
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems. It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu and Low, 2023).
Studying Toy Models for Arithmetic. Doing mechanistic interpretability on toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023). Quirke and Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on rare edge cases (e.g. “77778+22222=100000" where a “carry 1" cascades through 4 digits), highlighting the importance of understanding and testing all edge cases for trustworthiness.
Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized" into discrete chunks (quanta) (Michaud et al., 2023). Models must learn these quanta to reduce loss. Understanding a network reduces to enumerating its quanta. Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
@@ 3 Methodology
Transformer models may learn addition algorithms different from traditional human methods. We define an alternative, mathematically-equivalent framework for addition and demonstrate our model implements this approach.
@@ 3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. We define the first number as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. Figure 2 shows an illustrative example.
@@
First, we adopt the framework from Quirke and Barez (2024) for our model’s addition process. The “Simple Addition" sub-task $A_{n}.SA$, which naively calculates the sum of digit pairs, is defined as:
@@
When there is no carry bit from the previous digit $A_{n}=A_{n}.SA$. The “Simple Carry" sub-task $A_{n}.SC$ determines whether the addition creates a carry bit:
@@
While Quirke et al.’s model is capable of handling simple addition and carry bits generated directly from digit pair addition, it encounters difficulties with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
Consider “00144+00056=000210". Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position. A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110". The Quirke et al. 1-layer model could cascade carry bits across two digits, but not three or more.
@@ 3.2 Extending the Mathematical Framework
To answer “44444+55556=" with “100000", an accurate model must predict the first answer digit A5 as “1". To do so, an accurate model must implement a “carry one cascade" circuit, which combines the “carry one" information from all five digits. This is especially hard as the model predicts answer tokens from left to right.
We introduce a digit-level sub-task called TriCase that calculates the essential “carry one" information for a single pair of digits $D_{n}$ and $D^{\prime}_{n}$. TriCase has 3 possible outputs representing a definite carry one (ST10), a possible carry one depending on the results of other calculations (ST9), or definitely not a carry one (ST8):
$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10,\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9,\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8,\\ \end{cases}\end{split}$ (3)
To perform the “cascading carry one" calculation, we introduce a TriAdd sub-task. It handles the case where a possible carry one becomes a definite carry one because the next lower digit pair generated a carry one. TriAdd is defined as:
@@
The model can use $ST$ and $SV$ to accurately calculate A5 as 1 or 0 by using A5 = A4.SV = TriAdd(A4.ST, TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST)))). Note that in calculating A5, the model has also calculated an accurate carry bit for each answer digit. For example, the carry bit for A4 is A3.SV = TriAdd(A3.ST, TriAdd(A2.ST, TriAdd(A1.ST, A0.ST))).
With this framework, the model only needs the sub-tasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform addition (but some models also use the redundant $A_{n}$.SC sub-task). This framework, if implemented by a model, is sufficient for the model to perform n-digit addition accurately.
Figure 3 diagrams how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke et al. algorithm but contains an additional circuit (the shaded boxes) to calculate “cascading carry one" data to predict with 99.9999% accuracy.
Figure 3: To predict with 99.9999% accuracy, the addition algorithm first calculates “carry one" values $(A_{n}.ST)$, combining them into “cascading carry one" values $(A_{n}.SV).$ At the “+" token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0. The other answer digits are calculated by combining “base add" $(A_{n}.BA)$ and “carry one" $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
@@ 3.3 Techniques
In investigating the model circuits, we want to understand what each attention head or MLP layer is doing across each token position, and how it relates to our mathematical framework. Hence, we define a node as the computation done by an attention head or MLP layer for a given token position. To investigate what each node is doing, we use the following techniques:
@@
    Intervention Ablation. To find out how the model depends on the output of a node, we replace the output of that node with the vector that is the mean of all of its outputs across a batch and measure how that impacts downstream performance. We also use (automated, n-digit) intervention ablation tests, specific to each sub-task, to test for the expected sub-task behavior.
@@
    Attention Patterns. To find out what the model attends to at a node, we take the attention pattern at that token position and take the significant tokens attended to ($>0.01$ post softmax).
@@
    Principal Component Analysis (PCA). We use PCA to investigate the outputs of attention heads, especially where our framework suggests the head output may be tri-state or bi-state.
@@
    Question Complexity. We categorized questions by computational complexity (App. G). Addition categories (S0-S4) and subtraction categories (M0-M4) reflect the number of sequential digits a “carry one" or “borrow one" cascades through, respectively. We analyzed which nodes were necessary for correct predictions in each category.
@@ 4.1 Training a Five-Digit Addition Model
The Quirke et al. 5-digit 1-layer addition model achieved an accuracy of $\sim$ 99%. Our experiments suggested that a 2-layer, 3-head model was the smallest configuration capable of achieving 99.9999% accuracy (see App. D for alternatives tested). This configuration effectively doubled the computational power compared to the 1-layer model (see App. B for mode model configuration details). Moreover, a 2-layer model introduces the capability to “compose" the attention heads in novel ways, facilitating the implementation of more complex algorithms (Elhage et al., 2021b).
We trained a 5-digit, 2-layer, 3-head model, with a 14 token vocabulary ( 0, .., 9, +, -, =, *, / ), batch size of 64, learning rate of 0.00008 and weight decay of 0.1. Training used an infinite dataset enriched with rare edge cases. Loss was defined as the mean across all answer tokens of their negative log likelihood loss. After 30 thousand training batches, the model’s final training loss was $\sim 2.3\times 10^{-8}$. Testing showed this model has six nines accuracy. (More details in App. C and Tab. 6).
@@
Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact. Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5. For all addition and mixed models studied, before the "=" token, each node often calculates data used to predict multiple answer digits. After the "=" token, all nodes in a given token position are used to predict a single answer digit.

### v9

# Arithmetic in Transformers Explained

@@ Abstract
While recent work has shown transformers can learn addition, previous models exhibit poor prediction accuracy and are limited to small numbers. Furthermore, the relationship between single-task and multitask arithmetic capabilities remains unexplored. In this work, we analyze 44 autoregressive transformer models trained on addition, subtraction, or both. These include 16 addition-only models, 2 subtraction-only models, 8 “mixed” models trained to perform addition and subtraction, and 14 mixed models initialized with parameters from an addition-only model. The models span 5- to 15-digit questions, 2 to 4 attention heads, and 2 to 3 layers. We show that the addition models converge on a common logical algorithm, with most models achieving >99.999% prediction accuracy. We provide a detailed mechanistic explanation of how this algorithm is implemented within the network architecture. Subtraction-only models have lower accuracy. With the initialized mixed models, through parameter transfer experiments, we explore how multitask learning dynamics evolve, revealing that some features originally specialized for addition become polysemantic, serving both operations, and boosting subtraction accuracy. We explain the mixed algorithm mechanically. Finally, we introduce a reusable library of mechanistic interpretability tools to define, locate, and visualize these algorithmic circuits across multiple models.
Machine Learning, ICML
11footnotetext: Corresponding author: philipquirkenz@gmail.com
@@ 1 Introduction
Figure 1: An overview of our methodology: (1) We trained an accurate 5-digit integer addition model. (2) We reverse-engineered the model to find the algorithms that were implemented to perform addition. (3) We inserted the addition model into a new model, by copying the weights of the attention heads and MLPs (in brown) into the larger model during initialization. (4) We then train the new model on 80% subtraction and 20% addition questions. (5) We find that the resulting model predicts accurately and reuses the inserted addition circuits for both addition and subtraction questions.
@@
Transformer-based large language models (LLMs) are powerful (Barak et al., 2022) yet largely inscrutable due to their complex, nonlinear interactions in dense layers within high-dimensional spaces. Given this complexity, their deployment in critical settings (Zhang et al., 2022) highlights the need to understand their behavior. Hendrycks & Mazeika (2022) argue that making these models interpretable is key to their safe use.
Previous work by (Quirke & Barez, 2024) demonstrated that transformers can learn addition on small numbers, but their models had poor accuracy and were limited to a single operation. Understanding how transformers can overcome these limitations to perform more accurate mathematical computations across multiple operations is crucial for several reasons. First, it provides insights into how neural networks can learn and combine multiple precise algorithms, a key capability for more complex reasoning tasks. Second, studying how models transition from single-task to multitask competence while maintaining accuracy reveals important principles about knowledge transfer and representation sharing in neural networks.
Understanding how models arrive at their predictions is critical for diagnosing errors, enhancing safety, and gaining fundamental insights into modern neural networks. Mechanistic interpretability provides a pathway toward this understanding by identifying the “circuits” responsible for particular computations within a network’s layers and attention heads.
In this work, we use multidigit arithmetic as a revealing testbed for mechanistic interpretability. We train and analyze 44 autoregressive transformer models on addition, subtraction, or both, covering question lengths of 5 to 15 digits and varying in their number of layers (2–3) and attention heads (2–4). Our experiments encompass three categories of models: addition-only, subtraction-only, and “mixed” models trained to handle both operations. Through a systematic investigation of these networks, we show that addition-only models converge on a highly accurate, shared algorithm, yielding near-perfect performance. We mechanistically explain the algorithm. Subtraction-only models, in contrast, exhibit lower accuracy.
To explore how multitask learning reshapes these internal algorithms, we then conduct parameter transfer experiments using three distinct approaches, each involving the initialization of mixed models with parameters from addition-only networks. Interestingly, one of these approaches boosts the subtraction accuracy beyond what is achieved by mixed models trained from scratch, indicating that features specialized for addition can be effectively repurposed for subtraction. We provide a detailed mechanistic explanation of the mixed algorithm, illustrating how certain addition circuits become “polysemantic” under multitask learning, driving behavior for both operations.
A further contribution of our study is the development of a reusable library of mechanistic interpretability tools. This library enables researchers to define and search for algorithmic features, visualize network structures and feature interactions, and conduct cross-model comparisons efficiently.
Overall, our results demonstrate the power of mechanistic interpretability in shedding light on how transformers learn and reuse algorithmic circuits across related tasks, paving the way for deeper understanding and more reliable deployment of large language models in diverse domains.
@@ 2 Related Work
Investigating pre-trained LMs on Arithmetic. Even though basic arithmetic can be solved following a few simple rules, pre-trained LMs often struggle to solve simple math questions (Hendrycks et al., 2021). Causal mediation analysis (Stolfo et al., 2023) has been used to investigate how large pre-trained LMs like Pythia and GPT-J performed addition to solve word problems. It is also possible to improve a model’s arithmetic abilities used supervised fine tuning - including enriched training data (Liu & Low, 2023).
Studying Toy Models for Arithmetic. Applying mechanistic interpretability to toy transformers can help to better isolate clear, distinct circuits given the highly specific experimental setup for the model studied (Nanda et al., 2023). Quirke & Barez (2024) detailed a 1-layer, 3-head transformer model that performs 5-digit addition, showing it failed on edge cases (e.g. “77778+22222=100000” where a “carry 1” cascades through 4 digits), incorrectly predicting 1% of questions.
Many natural prediction problems decompose into a finite set of knowledge and skills that are “quantized” into discrete chunks (quanta) (Michaud et al., 2023). Models must learn these quanta to reduce loss. Understanding a network reduces to enumerating its quanta. Other studies (Schaeffer et al., 2023) prove useful ways to measure quanta in mathematical prediction problems.
@@ 3 Methodology
Transformer models learn addition algorithms different from traditional human methods. We define an alternative, mathematically-equivalent framework for addition and demonstrate our models implements this approach.
@@ 3.1 Mathematical Framework
Consider the task of adding two $n$-digit numbers together. We define the first number (or addend) as $D=\{D_{n-1},D_{n-2},\ldots,D_{0}\}$ and the second number as $D^{\prime}=\{D^{\prime}_{n-1},D^{\prime}_{n-2},\ldots,D^{\prime}_{0}\}$ and the answer (or sum) as $A=\{A_{n},A_{n-1},\ldots,A_{0}\}$. Figure 2 shows an illustrative example.
@@
We adopt the framework from Quirke & Barez (2024) for our model’s addition process, where they break down the task of multi-digit addition into an addition subtask and a carry subtask. Firstly, the “Simple Addition” subtask $A_{n}.SA$ naively calculates the sum of corresponding digit pairs in the addends:
@@
The “Simple Carry” subtask $A_{n}.SC$ then determines whether the addition between the two digits creates a carry bit:
@@
By combining these two subtasks, the model and framework from Quirke & Barez (2024) is capable of handling simple addition and carry bits generated directly from digit pair addition. However, it struggles with problems with ’cascading carry’ bits, where a carry bit from one digit position propagates to the next.
Consider “00144+00056=000210”. Adding 4+5 in the tens position doesn’t generate a carry bit directly, but a carry bit propagates from the ones position. A model only summing digit pairs and their direct carry bits would fail, producing an incorrect result like “00144+50006=000110”. The Quirke & Barez (2024) 1-layer model could cascade carry bits across two digits, but not three or more.
@@ 3.2 Extending the Mathematical Framework
To solve $n$-digit addition with high accuracy, we need to extend the basic framework to handle carry bits that cascade through multiple digits. Consider the problem ”44444+55556=100000”. For an autoregressive model to correctly predict the first answer digit as ”1”, it needs to track how the carry bit cascades from the rightmost digit all the way to the leftmost position in a single forward pass. This is particularly challenging for an autoregressive model that predicts answer tokens from left to right.
To solve this, we extend the mathematical framework with two key subtasks.
First, the digit-level subtask ”Sum Type” $A_{n}.ST$ analyzes each pair of digits and classifies their sum as either definitely causing a carry, possibly causing a carry, or definitely not causing a carry. This is implemented by a TriCase operation that maps each digit pair ($D_{n}$ and $D^{\prime}_{n}$), TriCase into one of three possible states:
$\begin{split}A_{n}.ST&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}ST10&\text{if }(D_{n}+D^{\prime}_{n})\geq 10\text{ (definite carry)}\\ ST9&\text{if }(D_{n}+D^{\prime}_{n})=9\text{ (possible carry)}\\ ST8&\text{if }(D_{n}+D^{\prime}_{n})\leq 8\text{ (no carry)}\\ \end{cases}\end{split}$ (3)
The key insight is that while some digit pairs definitely generate a carry (like 5+6), and others definitely don’t (like 2+3), pairs that sum to 9 (like 4+5) may or may not generate a carry depending on whether there’s an incoming carry bit from the next lower digit.
Then, we introduce the ”Sum Validation” subtask $A_{n}.SV$ that processes these carry states sequentially to determine whether any ”possible carries” become ”definite carries”. This subtask is implemented using a TriAdd operation that examines each digit’s sum type along with the sum type of the next lower digit:
@@
Revisiting the problem ”44444+55556=100000”, the model can accurately determine whether the first answer digit A5 should be 1 or 0 by examining all digit pairs in sequence. This chain of operations uses TriAdd to process the Sum Types from least to most significant digits:
$A1.SV$ $=\text{TriAdd}(A1.ST,A0.ST)$
$A2.SV$ $=\text{TriAdd}(A2.ST,A1.SV)$
$A3.SV$ $=\text{TriAdd}(A3.ST,A2.SV)$
$A4.SV$ $=\text{TriAdd}(A4.ST,A3.SV)$
$A5$ $=(A4.SV=ST10)$
Most importantly, by using this cascading calculation to calculating A5, the model also simulatenously calculates the accurate carry bits for each answer digit. The carry bits are simply the SV values we computed along the way—we already have A3.SV for A4’s carry bit, A2.SV for A3’s carry bit, and so on.
Within our extended framework, the model only requires the subtasks $A_{n}$.SA, $A_{n}.ST$ and $A_{n}.SV$ to accurately perform $n$-digit addition accurately. Some models may also use the redundant $A_{n}$.SC subtask.
Figure 3 shows how our model’s algorithm uses SA, SC, $ST$ and $SV$ to perform addition. This algorithm is like the 99% accurate Quirke & Barez (2024) algorithm, but contains an additional circuit (the shaded boxes) to calculate “cascading carry one” data to solve $n$-digit addition with >99.999% accuracy.
Figure 3: To predict with >99.999% accuracy, the addition algorithm first calculates “carry one” values $(A_{n}.ST)$, combining them into “cascading carry one” values $(A_{n}.SV).$ At the “+” token, $A_{n-1}.SV$ gives the first answer digit as 1 or 0. The other answer digits are calculated by combining “base add” $(A_{n}.BA)$ and “carry one” $(A_{n}.SC)$ calculations with the pre-calculated $A_{n}.SV$ values.
@@ 3.3 Techniques
To undetsand how the model implements arithmetic operations, we analyzed both attention heads and MLP layers at each token position. We define a node as either an attention head or MLP layer at a specific token position (e.g., the first attention head in the first layer processing the third digit). We employed four techniques to understand each node’s function:
@@
    Intervention Ablation. To determine the role of a node’s output, we replace its output with the mean vector of its outputs across a batch, then measured how this impacted downstream predictions. For each subtask, we developed automated intervention ablation tests across $n$-digits to verify whether a node implemented that subtask.
@@
    Attention Patterns. For each attention head node, we examined which input tokens received significant attention (defined as post-softmax attention score $>0.01$). This revealed which parts of the input each attention head considered relevant.
@@
    Principal Component Analysis (PCA). We applied PCA to node outputs, especially when investigating attention heads that our framework suggests might have descrete states (e.g. tri-state or bi-state outputs).
@@
    Question Complexity. We categorized questions by computational complexity (App. G). For addition, categories S0-S4 reflect how many sequential digits a “carry one” operation cascades through. For subtraction, categories M0-M4 reflect a similar cascading of ”borrow one” operations. We analyzed which nodes were necessary for correct predictions in each category.
@@ 4.1 Training a Five-Digit Addition Model
Previous work by Quirke & Barez (2024) achieved $\sim$ 99% accuracy on 5-digit addition using a 1-layer addition model. To improve upon this, we systematically explored different model architectures and found that a transformer with 2 layers and 3 attention heads per layer was the minimal configuration needed to achieve >99.999% accuracy (see App. D for our exploration of alternative architectures, and App. B for model configuration details.). This 2-layer architecture doubles the model’s computational capacity and enables cross-layer attention composition for more sophiscated algorithms (Elhage et al., 2021b).
We trained our model to perform 5-digit addition using an architecture with 2 layers and 3 attention heads per layer. The model’s vocabulary consisted of 14 tokens: the digits 0-9 plus the mathematical operators +, -, =, *, and /. We used a batch size of 64, learning rate of 0.00008, and weight decay of 0.1.
To ensure robust performance on edge cases, we trained on an infinite dataset that was deliberately enriched with rare numerical combinations. The loss function was defined as the mean negative log likelihood across all output tokens. After 30,000 training batches, the model achieved a final training loss of approximately $2.3\times 10^{-8}$, and subsequent testing confirmed accuracy exceeding 99.999%. For comprehensive training details and results, see App. C and Tab. 6.
Table 1: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact. Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5. For all addition and mixed models studied, before the ”=” token, each node often calculates data used to predict multiple answer digits. After the ”=” token, all nodes in a given token position are used to predict a single answer digit.